# Dependency Parsing is More Parameter-Efficient with Normalization

### v1

@@ B.1 Tagger
Table 5 reports the results for the best values of $h_{\psi}$ and $d_{\text{MLP}}$ (chosen as described in Section 4), ablating over $\phi\in\{\checkmark,\times\}$ and $\textbf{e}_{i}^{tag}\in\{\checkmark,\times\}$. Note that the first quarter of the table is equivalent to Table 3.
@@
B.2 Additional BiLSTM layers
[…]
As shown in Figure 5, performance is initially greater when normalizing scores, with $L_{\psi}~{}\in~{}\{0,1,2,3,4,5\}$. In this range, an increasing amount of layers reduces the amount of benefit obtained by using normalization. However, at $L_{\psi}>5$, for ADE, CoNNL04, and ERFGC, performance drops suddenly and rapidly when using biaffine score normalization. In the case of SciERC, the drop is more gentle, with the performance for the normalized and raw biaffine scores eventually converging, albeit with very high standard deviation at $L_{\psi}=10$.
As observed in Section 5, adding trainable layers seemingly allows the parameters to scale the variance of the scores to compensate for the lack of explicit normalization. As reported in Appendix B.6, a similar behavior can be observed with fully fine-tuned models, where normalization yields diminishing benefits the more parameters we train. This behavior points to a trade-off between the use of our technique and the amount of stacked BiLSTM layers, at a given amount of training steps.
B.3 MLP output dimension
As shown in Table 6, without normalization, increasing the output dimension of the two MLPs responsible for projecting edge representation projections leads to a decrease in performance. This is in contrast with the behavior observed when using normalization, where performance is essentially independent from the output dimension. This is in line with our claim that higher variance in the edge scores causes lower performance. Normalizing the scores assuages the variance, which makes performance stable even at high output dimensions. Once again, these results show the relationship between score variance and performance, with equivalent performance being achievable with fewer parameters.
Table 6: Performance in terms of F1-measure when ablating over different output dimension for the parser’s MLPs ($\phi=\checkmark$, $\textbf{e}_{i}^{tag}=\checkmark$). Best in bold.
@@ B.3 MLP output dimension
$(L_{\psi},h_{\psi})=$ (3, 200) (3, 400) (3, 300) (3, 400)
$1$ 100 0.662 ±$\pm 0.027$ 0.579 ±$\pm 0.030$ 0.302 ±$\pm 0.018$ 0.709 ±$\pm 0.008$
300 0.658 ±$\pm 0.018$ 0.562 ±$\pm 0.021$ 0.299 ±$\pm 0.023$ 0.705 ±$\pm 0.011$
500 0.657 ±$\pm 0.018$ 0.566 ±$\pm 0.019$ 0.281 ±$\pm 0.016$ 0.701 ±$\pm 0.009$
$\frac{1}{\sqrt{d}}$ 100 0.686 ±$\pm 0.025$ 0.586 ±$\pm 0.013$ 0.318 ±$\pm 0.017$ 0.712 ±$\pm 0.008$
300 0.680 ±$\pm 0.018$ 0.602 ±$\pm 0.017$ 0.320 ±$\pm 0.013$ 0.708 ±$\pm 0.008$
500 0.685 ±$\pm 0.019$ 0.600 ±$\pm 0.015$ 0.311 ±$\pm 0.009$ 0.707 ±$\pm 0.004$
B.4 BiLSTM hidden size
As Table 7 shows, the hidden size of the BiLSTMs does not have a visible effect on performance. This is especially the case when applying score normalization, which produces smaller standard deviations. As a result, not only do models perform better with biaffine score normalization, but performance is also less dependent on the hidden size of the BiLSTM encoders. This supports our claim that score normalization is a useful technique to obtain the same performance with lower parameter count, since the models tend to perform comparably, despite the lower BiLSTM hidden sizes.
@@ B.4 BiLSTM hidden size
$(L_{\psi},d_{\text{MLP}})=$ (3, 100) (3, 300) (3, 300) (3, 300)
$1$ 100 0.682 ±$\pm 0.021$ 0.580 ±$\pm 0.031$ 0.291 ±$\pm 0.021$ 0.693 ±$\pm 0.011$
200 0.662 ±$\pm 0.027$ 0.582 ±$\pm 0.006$ 0.286 ±$\pm 0.032$ 0.703 ±$\pm 0.007$
300 0.674 ±$\pm 0.029$ 0.585 ±$\pm 0.026$ 0.299 ±$\pm 0.023$ 0.703 ±$\pm 0.006$
400 0.663 ±$\pm 0.032$ 0.562 ±$\pm 0.021$ 0.289 ±$\pm 0.038$ 0.705 ±$\pm 0.011$
$\frac{1}{\sqrt{d}}$ 100 0.685 ±$\pm 0.020$ 0.600 ±$\pm 0.018$ 0.302 ±$\pm 0.013$ 0.698 ±$\pm 0.008$
200 0.686 ±$\pm 0.025$ 0.610 ±$\pm 0.018$ 0.314 ±$\pm 0.019$ 0.702 ±$\pm 0.008$
300 0.678 ±$\pm 0.012$ 0.599 ±$\pm 0.012$ 0.320 ±$\pm 0.013$ 0.711 ±$\pm 0.005$
400 0.674 ±$\pm 0.011$ 0.602 ±$\pm 0.017$ 0.306 ±$\pm 0.016$ 0.708 ±$\pm 0.008$
B.5 LayerNorm and parameter initialization
In this section, we analyze the effects of using a Xavier normal distribution ($I_{par}=\mathcal{N}$) [^4] https://docs.pytorch.org/docs/stable/nn.init.html#torch.nn.init.xavier_normal_ for the weights of the two projections $\text{MLP}^{(edge-head)}$ and $\text{MLP}^{(edge-dept)}$, along with the edge biaffine layer $f^{(edge)}(\,\cdot\ ;W_{e})$. We also apply a LayerNorm function $\texttt{LN}_{\psi}$ [2] for each of the BiLSTM layers.
As reported in Table 8, on average the best results are obtained with the base setting, i.e. with $I_{par}=\mathcal{U}$ and $\texttt{LN}_{\psi}=\times$. Using LayerNorm can provide a performance boost for some datasets, especially with uniform initialization. However, it makes performance unstable for some. As a matter of fact, when using LayerNorm with three BiLSTM layers on CoNLL04 and SciERC, the model often breaks and is not able to converge. Based on average performance, then, the best models are obtained by scaling biaffine scores ($a=1/\sqrt{d}$) and initializing the parser with a uniform weight distribution, without any LayerNorm in between the stacked BiLSTM layers.
Table 8: Ablation over the best hyperparameters ($h_{\psi},d_{\text{MLP}}$) and $\phi=\checkmark$ and $\textbf{e}_{i}^{tag}=\checkmark$, using LayerNorm layers $\texttt{LN}_{\psi}\in\{\checkmark,\times\}$ and Xavier uniform/normal initialization $I_{par}\in\{\mathcal{U},\mathcal{N}\}$ for the parser. The first quarter of the table is equivalent to Table 3. Best in bold, second-best underlined.
$I_{par}$ $\texttt{LN}_{\psi}$ $a$ $L_{\psi}$ ADE CoNLL04 SciERC ERFGC Mean
@@ B.5 LayerNorm and parameter initialization
$\mathcal{U}$square $1$ 0 0.541 ±$\pm 0.021$ 0.399 ±$\pm 0.024$ 0.147 ±$\pm 0.049$ 0.548 ±$\pm 0.010$ 0.515
$\square$ 3 0.662 ±$\pm 0.027$ 0.562 ±$\pm 0.021$ 0.299 ±$\pm 0.023$ 0.705 ±$\pm 0.011$
$\square$ 1 0.657 ±$\pm 0.011$ 0.556 ±$\pm 0.021$ 0.282 ±$\pm 0.009$ 0.676 ±$\pm 0.010$
$\square$ 2 0.667 ±$\pm 0.011$ 0.573 ±$\pm 0.025$ 0.273 ±$\pm 0.010$ 0.694 ±$\pm 0.010$
$\square$ $\frac{1}{\sqrt{d}}$ 0 0.567 ±$\pm 0.014$ 0.438 ±$\pm 0.033$ 0.181 ±$\pm 0.027$ 0.612 ±$\pm 0.008$ 0.541
$\square$ 1 0.668 ±$\pm 0.017$ 0.597 ±$\pm 0.015$ 0.299 ±$\pm 0.019$ 0.692 ±$\pm 0.009$
$\square$ 2 0.676 ±$\pm 0.019$ 0.596 ±$\pm 0.014$ 0.312 ±$\pm 0.011$ 0.699 ±$\pm 0.009$
$\square$ 3 0.686 ±$\pm 0.025$ 0.602 ±$\pm 0.017$ 0.320 ±$\pm 0.013$ 0.708 ±$\pm 0.008$
$\blacksquare$ $1$ 0 0.545 ±$\pm 0.018$ 0.399 ±$\pm 0.024$ 0.151 ±$\pm 0.014$ 0.548 ±$\pm 0.010$ 0.468
$\blacksquare$ 1 0.680 ±$\pm 0.014$ 0.554 ±$\pm 0.042$ 0.265 ±$\pm 0.028$ 0.686 ±$\pm 0.007$
$\blacksquare$ 2 0.679 ±$\pm 0.011$ 0.578 ±$\pm 0.033$ 0.260 ±$\pm 0.020$ 0.715 ±$\pm 0.012$
$\blacksquare$ 3 0.680 ±$\pm 0.013$ 0.117 ±$\pm 0.261$ 0.060 ±$\pm 0.133$ 0.569 ±$\pm 0.318$
$\blacksquare$ $\frac{1}{\sqrt{d}}$ 0 0.567 ±$\pm 0.014$ 0.433 ±$\pm 0.029$ 0.181 ±$\pm 0.027$ 0.614 ±$\pm 0.004$ 0.503
$\blacksquare$ 1 0.664 ±$\pm 0.017$ 0.564 ±$\pm 0.037$ 0.282 ±$\pm 0.029$ 0.686 ±$\pm 0.004$
$\blacksquare$ 2 0.697 ±$\pm 0.022$ 0.623 ±$\pm 0.019$ 0.300 ±$\pm 0.042$ 0.702 ±$\pm 0.011$
$\blacksquare$ 3 0.675 ±$\pm 0.019$ 0.239 ±$\pm 0.327$ 0.111 ±$\pm 0.152$ 0.703 ±$\pm 0.006$
$\mathcal{N}$square $1$ 0 0.545 ±$\pm 0.017$ 0.415 ±$\pm 0.014$ 0.155 ±$\pm 0.019$ 0.558 ±$\pm 0.009$ 0.520
$\square$ 1 0.667 ±$\pm 0.014$ 0.543 ±$\pm 0.017$ 0.275 ±$\pm 0.020$ 0.680 ±$\pm 0.015$
$\square$ 2 0.674 ±$\pm 0.025$ 0.576 ±$\pm 0.023$ 0.272 ±$\pm 0.014$ 0.699 ±$\pm 0.006$
$\square$ 3 0.672 ±$\pm 0.028$ 0.580 ±$\pm 0.022$ 0.297 ±$\pm 0.019$ 0.705 ±$\pm 0.006$
$\square$ $\frac{1}{\sqrt{d}}$ 0 0.570 ±$\pm 0.013$ 0.454 ±$\pm 0.006$ 0.181 ±$\pm 0.023$ 0.613 ±$\pm 0.007$ 0.538
$\square$ 1 0.671 ±$\pm 0.015$ 0.578 ±$\pm 0.031$ 0.299 ±$\pm 0.030$ 0.685 ±$\pm 0.010$
$\square$ 2 0.671 ±$\pm 0.031$ 0.593 ±$\pm 0.016$ 0.301 ±$\pm 0.019$ 0.704 ±$\pm 0.007$
$\square$ 3 0.681 ±$\pm 0.024$ 0.590 ±$\pm 0.014$ 0.315 ±$\pm 0.009$ 0.702 ±$\pm 0.011$
$\blacksquare$ $1$ 0 0.545 ±$\pm 0.017$ 0.415 ±$\pm 0.014$ 0.155 ±$\pm 0.019$ 0.558 ±$\pm 0.009$ 0.469
$\blacksquare$ 1 0.679 ±$\pm 0.012$ 0.582 ±$\pm 0.012$ 0.259 ±$\pm 0.010$ 0.680 ±$\pm 0.016$
$\blacksquare$ 2 0.668 ±$\pm 0.015$ 0.580 ±$\pm 0.041$ 0.284 ±$\pm 0.024$ 0.711 ±$\pm 0.005$
$\blacksquare$ 3 0.679 ±$\pm 0.018$ $0$$\pm 0.000$ $0$$\pm 0.000$ 0.711 ±$\pm 0.009$
$\blacksquare$ $\frac{1}{\sqrt{d}}$ 0 0.570 ±$\pm 0.013$ 0.454 ±$\pm 0.006$ 0.181 ±$\pm 0.023$ 0.613 ±$\pm 0.007$ 0.512
$\blacksquare$ 1 0.675 ±$\pm 0.019$ 0.578 ±$\pm 0.022$ 0.280 ±$\pm 0.022$ 0.682 ±$\pm 0.006$
$\blacksquare$ 2 0.687 ±$\pm 0.021$ 0.609 ±$\pm 0.012$ 0.308 ±$\pm 0.012$ 0.702 ±$\pm 0.011$
$\blacksquare$ 3 0.678 ±$\pm 0.009$ 0.476 ±$\pm 0.266$ $0$$\pm 0.000$ 0.701 ±$\pm 0.006$
B.6 Full fine-tuning and sample efficiency

### v2

@@ D.1 Tagger
Table 8 reports the results for the best values of $h_{\psi}$ and $d_{\text{MLP}}$ (chosen as described in Section 4), ablating over $\phi\in\{\checkmark,\times\}$ and $\textbf{e}_{i}^{tag}\in\{\checkmark,\times\}$. The first quarter of the table is equivalent to Table 3.
@@
D.2 Additional BiLSTM layers
Figure 6: Micro-F1 (SemDP) and LAS (SynDP) vs $N\in\{0,\dots,10\}$ at 20kk training steps. Best per-dataset hyperparameters ($h_{\psi}$, $d_{\text{MLP}}$) with $\phi=\checkmark$ and $\mathbf{e}^{tag}_{i}=\checkmark$. Red = norm; blue = raw.
Figure 6 depicts performance against the number of BiLSTM layers $N$ at 20kk training steps. Compared with the main setting’s 2kk steps, we increase the amount of training because lower amounts resulted in very high F1 standard deviations with as few as $N=6$ layers in preliminary experiments. With a shallow stack of BiLSTM layers ($N<6$), performance is greater when normalizing scores. However, at deeper depths performance becomes very unstable for CoNLL04, ERFGC, and enEWT. Nonetheless, it is evident that score normalization is beneficial, with deeper stacks of BiLSTM layers reducing the effect of normalization, supporting our main-setting claims and experiments.
As observed in Section 5, adding trainable layers seemingly allows the parameters to scale the variance of the scores to compensate for the lack of explicit normalization. As reported in Appendix D.6, a similar behavior can be observed with fully fine-tuned models, where normalization yields diminishing benefits the more parameters we train. This behavior points to a trade-off between the use of our technique and the amount of stacked BiLSTM layers, on a given dataset and at a given amount of training steps.
D.3 MLP output dimension
As shown in Table 9, without normalization, increasing the output dimension of the two MLPs responsible for projecting edge representations leads to a decrease in performance. This is in contrast with the behavior observed when using normalization, where performance is essentially independent from the output dimension. This is in line with our claim that higher variance in the edge scores causes lower performance. Normalizing the scores reduces the variance, which makes performance stable even at high output dimensions. Once again, these results show the relationship between score variance and performance, with equivalent performance being achievable with fewer parameters.
Table 9: Performance in terms of F1-measure when ablating over different output dimensions for the parser’s MLPs ($\phi=\checkmark$, $\textbf{e}_{i}^{tag}=\checkmark$). Best in bold.
@@ D.3 MLP output dimension
$(N,h_{\psi})=$ (3, 200) (3, 400) (3, 300) (3, 400)
11 100 0.662 $\pm 0.027$ 0.579 $\pm 0.030$ 0.302 $\pm 0.018$ 0.709 $\pm 0.008$
300 0.658 $\pm 0.018$ 0.562 $\pm 0.021$ 0.299 $\pm 0.023$ 0.705 $\pm 0.011$
500 0.657 $\pm 0.018$ 0.566 $\pm 0.019$ 0.281 $\pm 0.016$ 0.701 $\pm 0.009$
$\frac{1}{\sqrt{d}}$ 100 0.686 $\pm 0.025$ 0.586 $\pm 0.013$ 0.318 $\pm 0.017$ 0.712 $\pm 0.008$
300 0.680 $\pm 0.018$ 0.602 $\pm 0.017$ 0.320 $\pm 0.013$ 0.708 $\pm 0.008$
500 0.685 $\pm 0.019$ 0.600 $\pm 0.015$ 0.311 $\pm 0.009$ 0.707 $\pm 0.004$
D.4 BiLSTM hidden size
As Table 10 shows, the hidden size of the BiLSTMs does not have a visible effect on performance. This is especially the case when applying score normalization, which produces smaller standard deviations. As a result, not only do models perform better with biaffine score normalization, but performance is also less dependent on the hidden size of the BiLSTM encoders. This supports our claim that score normalization is a useful technique to obtain the same performance with lower parameter counts, since the models tend to perform comparably, despite lower BiLSTM hidden sizes.
@@ D.4 BiLSTM hidden size
$(N,d_{\text{MLP}})=$ (3, 100) (3, 300) (3, 300) (3, 300)
11 100 0.682 $\pm 0.021$ 0.580 $\pm 0.031$ 0.291 $\pm 0.021$ 0.693 $\pm 0.011$
200 0.662 $\pm 0.027$ 0.582 $\pm 0.006$ 0.286 $\pm 0.032$ 0.703 $\pm 0.007$
300 0.674 $\pm 0.029$ 0.585 $\pm 0.026$ 0.299 $\pm 0.023$ 0.703 $\pm 0.006$
400 0.663 $\pm 0.032$ 0.562 $\pm 0.021$ 0.289 $\pm 0.038$ 0.705 $\pm 0.011$
$\frac{1}{\sqrt{d}}$ 100 0.685 $\pm 0.020$ 0.600 $\pm 0.018$ 0.302 $\pm 0.013$ 0.698 $\pm 0.008$
200 0.686 $\pm 0.025$ 0.610 $\pm 0.018$ 0.314 $\pm 0.019$ 0.702 $\pm 0.008$
300 0.678 $\pm 0.012$ 0.599 $\pm 0.012$ 0.320 $\pm 0.013$ 0.711 $\pm 0.005$
400 0.674 $\pm 0.011$ 0.602 $\pm 0.017$ 0.306 $\pm 0.016$ 0.708 $\pm 0.008$
D.5 LayerNorm and parameter initialization
In this section, we analyze the effects of using a Xavier normal initialization ($I_{par}\sim\mathcal{N}$) [^7] https://docs.pytorch.org/docs/stable/$n$.init.html#torch.nn.init.xavier_normal_ for the weights of the two projections $\text{MLP}^{(edge-head)}$ and $\text{MLP}^{(edge-dept)}$, along with the edge biaffine layer $f^{(edge)}(\,\cdot\ ;W_{e})$. We also apply a LayerNorm function $\texttt{LN}_{\psi}$ [2] for each of the BiLSTM layers.
As reported in Table 11, on average the best results are obtained with the base setting, i.e., with $I_{par}\sim\mathcal{U}$ and $\texttt{LN}_{\psi}=\times$. Using LayerNorm can provide a performance boost for some datasets, especially with uniform initialization. However, it makes performance unstable for some. As a matter of fact, when using LayerNorm with three BiLSTM layers on CoNLL04 and SciERC, the model often breaks and is not able to converge. Based on average performance, then, the best models are obtained by scaling biaffine scores ($a=1/\sqrt{d}$) and initializing the parser with a uniform weight distribution, without any LayerNorm in between the stacked BiLSTM layers.
Table 11: Ablation over the best hyperparameters ($h_{\psi},d_{\text{MLP}}$) and $\phi=\checkmark$ and $\textbf{e}_{i}^{tag}=\checkmark$, using LayerNorm layers $\texttt{LN}_{\psi}\in\{\checkmark,\times\}$ and Xavier uniform/normal initialization $I_{par}\sim\{\mathcal{U},\mathcal{N}\}$ for the parser. The first quarter of the table is equivalent to Table 3. Best in bold, second-best underlined.
$I_{par}$ $\texttt{LN}_{\psi}$ $a$ NN ADE CoNLL04 SciERC ERFGC Mean
@@ D.5 LayerNorm and parameter initialization
$\mathcal{U}$ $\square$ 11 0 0.541 $\pm 0.021$ 0.399 $\pm 0.024$ 0.147 $\pm 0.049$ 0.548 $\pm 0.010$ 0.515
$\square$ 3 0.662 $\pm 0.027$ 0.562 $\pm 0.021$ 0.299 $\pm 0.023$ 0.705 $\pm 0.011$
$\square$ 1 0.657 $\pm 0.011$ 0.556 $\pm 0.021$ 0.282 $\pm 0.009$ 0.676 $\pm 0.010$
$\square$ 2 0.667 $\pm 0.011$ 0.573 $\pm 0.025$ 0.273 $\pm 0.010$ 0.694 $\pm 0.010$
$\square$ $\frac{1}{\sqrt{d}}$ 0 0.567 $\pm 0.014$ 0.438 $\pm 0.033$ 0.181 $\pm 0.027$ 0.612 $\pm 0.008$ 0.541
$\square$ 1 0.668 $\pm 0.017$ 0.597 $\pm 0.015$ 0.299 $\pm 0.019$ 0.692 $\pm 0.009$
$\square$ 2 0.676 $\pm 0.019$ 0.596 $\pm 0.014$ 0.312 $\pm 0.011$ 0.699 $\pm 0.009$
$\square$ 3 0.686 $\pm 0.025$ 0.602 $\pm 0.017$ 0.320 $\pm 0.013$ 0.708 $\pm 0.008$
$\blacksquare$ 11 0 0.545 $\pm 0.018$ 0.399 $\pm 0.024$ 0.151 $\pm 0.014$ 0.548 $\pm 0.010$ 0.468
$\blacksquare$ 1 0.680 $\pm 0.014$ 0.554 $\pm 0.042$ 0.265 $\pm 0.028$ 0.686 $\pm 0.007$
$\blacksquare$ 2 0.679 $\pm 0.011$ 0.578 $\pm 0.033$ 0.260 $\pm 0.020$ 0.715 $\pm 0.012$
$\blacksquare$ 3 0.680 $\pm 0.013$ 0.117 $\pm 0.261$ 0.060 $\pm 0.133$ 0.569 $\pm 0.318$
$\blacksquare$ $\frac{1}{\sqrt{d}}$ 0 0.567 $\pm 0.014$ 0.433 $\pm 0.029$ 0.181 $\pm 0.027$ 0.614 $\pm 0.004$ 0.503
$\blacksquare$ 1 0.664 $\pm 0.017$ 0.564 $\pm 0.037$ 0.282 $\pm 0.029$ 0.686 $\pm 0.004$
$\blacksquare$ 2 0.697 $\pm 0.022$ 0.623 $\pm 0.019$ 0.300 $\pm 0.042$ 0.702 $\pm 0.011$
$\blacksquare$ 3 0.675 $\pm 0.019$ 0.239 $\pm 0.327$ 0.111 $\pm 0.152$ 0.703 $\pm 0.006$
$\mathcal{N}$ $\square$ 11 0 0.545 $\pm 0.017$ 0.415 $\pm 0.014$ 0.155 $\pm 0.019$ 0.558 $\pm 0.009$ 0.520
$\square$ 1 0.667 $\pm 0.014$ 0.543 $\pm 0.017$ 0.275 $\pm 0.020$ 0.680 $\pm 0.015$
$\square$ 2 0.674 $\pm 0.025$ 0.576 $\pm 0.023$ 0.272 $\pm 0.014$ 0.699 $\pm 0.006$
$\square$ 3 0.672 $\pm 0.028$ 0.580 $\pm 0.022$ 0.297 $\pm 0.019$ 0.705 $\pm 0.006$
$\square$ $\frac{1}{\sqrt{d}}$ 0 0.570 $\pm 0.013$ 0.454 $\pm 0.006$ 0.181 $\pm 0.023$ 0.613 $\pm 0.007$ 0.538
$\square$ 1 0.671 $\pm 0.015$ 0.578 $\pm 0.031$ 0.299 $\pm 0.030$ 0.685 $\pm 0.010$
$\square$ 2 0.671 $\pm 0.031$ 0.593 $\pm 0.016$ 0.301 $\pm 0.019$ 0.704 $\pm 0.007$
$\square$ 3 0.681 $\pm 0.024$ 0.590 $\pm 0.014$ 0.315 $\pm 0.009$ 0.702 $\pm 0.011$
$\blacksquare$ 11 0 0.545 $\pm 0.017$ 0.415 $\pm 0.014$ 0.155 $\pm 0.019$ 0.558 $\pm 0.009$ 0.469
$\blacksquare$ 1 0.679 $\pm 0.012$ 0.582 $\pm 0.012$ 0.259 $\pm 0.010$ 0.680 $\pm 0.016$
$\blacksquare$ 2 0.668 $\pm 0.015$ 0.580 $\pm 0.041$ 0.284 $\pm 0.024$ 0.711 $\pm 0.005$
$\blacksquare$ 3 0.679 $\pm 0.018$ 0.000 $\pm 0.000$ 0.000 $\pm 0.000$ 0.711 $\pm 0.009$
$\blacksquare$ $\frac{1}{\sqrt{d}}$ 0 0.570 $\pm 0.013$ 0.454 $\pm 0.006$ 0.181 $\pm 0.023$ 0.613 $\pm 0.007$ 0.512
$\blacksquare$ 1 0.675 $\pm 0.019$ 0.578 $\pm 0.022$ 0.280 $\pm 0.022$ 0.682 $\pm 0.006$
$\blacksquare$ 2 0.687 $\pm 0.021$ 0.609 $\pm 0.012$ 0.308 $\pm 0.012$ 0.702 $\pm 0.011$
$\blacksquare$ 3 0.678 $\pm 0.009$ 0.476 $\pm 0.266$ 0.000 $\pm 0.000$ 0.701 $\pm 0.006$
D.6 Full fine-tuning and sample efficiency